as of Friday, 03-Dec-2021 07:10:31 AST

Go to Most Recent Entry

lswlwray@naic.eduMonday, April 26, 2004 13:11:32Receiver installed on April 21, 2004. First light on April 24.
lswlwray@naic.eduMonday, April 26, 2004 13:14:221200 - 1500 MHz bandpass filters installed on beam 0, pol A and B, between the cryogenic amplifier and the downconverter box.
lswlwray@naic.eduMonday 3, 2004 08:35:41low pass filters (cut off 1550 MHz) installed on all beams pol A only.
ansanolla@naic.eduThursday, May 6, 2004 11:38:00There is a He leak in the ALFA 1 (ED014) cryocomp . It is leaking around 10 to 15 p/day. Suspect the elbows next to the rf cabinets.
lswlwray@naic.eduMonday, May 10, 2004 08:34:35low pass filters removed from beam 0 ch A and beam 1 ch A, and left open in order to do a test at 1667 MHz.
ansanolla@naic.eduTuesday, May 11, 2004 13:29:17JVR swap cryo line K to cryo line 12 at the ALFA dewar.
lswlwray@naic.eduThursday, May 13, 2004 13:29:17Low pass filters put back on all beams both channels.
ansanolla@naic.eduMonday, May 17, 2004 13:26:52ALFA protected due to 430 mhz dual beam run. JVR mesh used (paper clips are used to attach it to Dewar).
ansanolla@naic.eduThursday, May 27, 2004 13:39:36Now there is an ALFA if monitor in rack 2. Now you can monitor manually the 14 if channels from the receiver , but remember to set up the ALFA first by typing "ifsetup alfa" at the vw% prompt.
lswlwray@naic.eduMonday, June 7, 2004 13:32:03Cryogenics: ALFA 1 compressor ED014 lost 80 pounds in 4 days, on Monday AM it was at 205/90. Checking for leaks...
lswlwray@naic.eduTuesday, June 15, 2004 08:29:28June 14: ALFA protected for 430 dual beam run. Cover put on. At advice of Ganesh, kept the control box on "remote" (not "local") and the display on "sleep". No changes to the motor control box.
lswlwray@naic.eduThursday, June 24, 2004 15:54:01See entry for LBW. Testing level of 2380 birdie injected before the transmitter horn seen after the ALFA's LNAs. Tested levels with and without the original RFI protection cover and Vives' new ring-clamp mesh hat. Results to be posted soon...
ansanolla@naic.eduFriday, July 2, 2004 11:24:24System down due to power shut down (shorted main power line at the rotary joint)
ansanolla@naic.eduTuesday, July 6, 2004 10:51:44System cool down started this morning.The vacuum chamber # 1 had a vacuum of 5 un's after 48 hrs. , chamber # 2 was 2 un's.
ansanolla@naic.eduWednesday, July 7, 2004 14:03:04System back to normal. It was out since 7/2/2004 due to a shorted power line at the rotary joint that warmed up all the systems.
ansanolla@naic.eduMonday, July 12, 2004 13:34:40ALFA set for 1665 mhz operation (18 cm). 18 cm rf filters on it.
ansanolla@naic.eduWednesday, July 14, 2004 10:51:52Channel 4B (#26) rf filter (18 cm,1665 mhz) on ALFA causing unstability on it. It's the main suspect.It was replaced by #30.
lswlwray@naic.eduThursday, July 15, 2004 16:27:27During a power dip today, the system was in remote and the amplifiers were on. When the power came back ~10 seconds later, the amplifiers were off.
lswllwray@naic.eduThursday, July 15, 2004 16:30:22looking for cryo He leak on system , line 11 replaced by 19 in order to find out the leak.Line 11 with monitor meters to check it.
ansanolla@naic.eduWednesday, July 21, 2004 08:20:08cryoline 19 was replaced with line 20 ,19 has a leaky connector solder at the end of the section coming from cryocompressors at the cable wrap , also it had a big leak at the elbow next to the rf cabinet that Vives fixed.ALFA Dewar cryotemps increased from 12/68 deg to 16/88 deg.Ganesh sugested to pump the Dewar.
ansanolla@naic.eduSunday, August 1, 2004 08:18:19The 18 cm (1665 mhz)rf filters replaced by the 21 cm. ones (1420 mhz) at AST 07:30. System checked by Steve.
ansanolla@naic.eduWednesday, August 4, 2004 11:43:17ALFA cryocomp # 2 (ED 015) was replace due to a vibration on it and funny noises. ED 008 on it's place.
ansanolla@naic.eduThursday, August 5, 2004 07:17:49Warming up system to purge both cryolines.Both cryocomps & cryoheads shut off.
lswlwray@naic.eduThursday, August 5, 2004 11:54:22The 70K stage cryo temperatures have been increasing from ~70 to >100K. Decided to purge the system using liquid nitrogen and the absolute filter.
ansanolla@naic.eduTuesday, August 10, 2004 14:27:01Cryocomp ED 008 (ALFA 1) is leaking He at a rate of 10 psi/day. On Friday these cryocomp will be swap with cryocomp ED 015.Also r.f. amp 0A 2nd stage is bad (shorted out).
ansanolla@naic.eduWednesday, August 11, 2004 15:04:03RF amplifier 0A 2nd stage shorted out.
ansanolla@naic.eduFriday, August 13, 2004 14:44:04ALFA cryocomp 1 (ED 008) was replaced by ED 015 , it had a 10 PSI leak on it (main suspect). Also the protective cover was set on it.
lswlwray@naic.eduThursday, August 19, 2004 14:10:41August 17: Amplifier that shorted out on Aug 11 (polA beam 0) is now OK. No known reason why it started working.
ansanolla@naic.eduWednesday, September 8, 2004 11:34:31Channel 00 converter (EC box) has an oscillation on it.Was replaced by the spare one.
ansanolla@naic.eduFriday, September 10, 2004 15:00:27ALFA cryo system turned off 'till next week.
ansanolla@naic.eduTuesday, September 14, 2004 14:44:42Cool down started at AST 13:30 today.Vacuum in 10 un's at B , 5 un's at A.
ansanolla@naic.eduWednesday, September 22, 2004 12:59:57ALFA UPS installed , UPS is at 50% load capacity.
ansanolla@naic.eduThursday, September 30, 2004 13:36:21ALFA was calibrated again.
ansanolla@naic.eduTuesday, November 30, 2004 13:03:53ALFA cryo comp # 1 (ED 015) stop working yesterday evening , Bill and Ganesh tried to solve the problem , but no way , it didn't started. So today during morning time Coco and Carlos went upstairs and found that the time delay relay was bad , they replaced it and cryo comp was on duty again. By the way Willy and Luis Murray modified the ED 007 cryo comp that was as back up but without the gimble bracket.
ansanolla@naic.eduThursday, December 2, 2004 10:14:44Ths SB xmitter was run without covering the ALFA.Ganesh is testing the system performance , because amps are acting strange.
ansanolla@naic.eduTuesday, January 11, 2005 13:45:34Cryocomp ED015 (ALFA #1) stop running. Coco check it and found that there was a loose relay on it.
ansanolla@naic.eduMonday, January 24, 2005 11:42:02ALFA cryocomp ED014 has a He leak of 10 psi/day.We try to localized but didn't find it.As far as we know cryo line #20 was swap with cryoline #11.
lswlwray@naic.eduFriday, January 28, 2005 13:56:26ALFA cryogenics have been turned off as of 1:30pm today. Motors are off and amplifiers are off also. The system will warm up until Monday Jan 31. Cool-down is planned for Tuesday Feb 1, should be cold by Wednesday.
lswlwray@naic.eduMonday, January 31, 2005 10:04:04Intermittent short in amplifier 1B stage 3 was discovered. Seems it appeared after a SBN radar run on Saturday night. Decided to warm up the dewar in hopes of eliminating the short.
ansanolla@naic.eduTuesday, February 1, 2005 12:30:37System been cooled down.
lswlwray@naic.eduTuesday, February 1, 2005 14:01:14Problems with reading dewar temperatures during the warm-up alerted our attention. The temperatures read anywhere from 220K to 280K, but the dewar is warm. The online monitor and the upstairs monitor both read the same. We swapped out both F30 temperature cards with the spares and the problem remained. Then we tried swapping the Analog Mux board (F82). The problem remains.
ansanolla@naic.eduWednesday, February 2, 2005 14:43:20System cooled down again.Short at amp 1B stage 3 gone.But something is wrong with cryotemp reading , too low at the 16 stage.
ansanolla@naic.eduMonday, March 7, 2005 12:59:27RF filters ( 21 Cm)removed in order to install them in the ALFA filter bank.
lswlwray@naic.eduWednesday, March 16, 2005 16:09:43Carlos, Cheo and Nolla installed the ALFA filter bank. It has a two position switch for each channel that switches manually (for now) between a 1550 Low-Pass and a 1660 (130 MHZ bandwidth) Bandwidth. Also, inside the filter bank, filter #26 was swapped for filter #31. There were reported instability problems with #26 in the past.
ansanolla@naic.eduThursday, March 31, 2005 14:56:54ALFA filter bank in the 18 cm mode (1665 Mhz).
ansanolla@naic.eduFriday, April 1, 2005 13:18:07Lambda P.S. mounted at the filter power supply.System still tuned to 18 cm (1665 Mhz).
ansanolla@naic.eduMonday, April 4, 2005 11:38:54ALFA 2 (014) cryo comp replaced with the SBH one (009)due that the return preassure was to high (135 psi).SBH down for a while.
ansanolla@naic.eduFriday, April 8, 2005 11:40:38On 4/5/2005 an 11 db gain Cougar amp was added to beam 02 (both channels)in order to increase its level for the 18 Cm side (1665 Mhz). Also a 3 db pad was added downstairs at the receiver Optical fiber at each channel.
ansanolla@naic.eduTuesday, April 19, 2005 14:11:22ALFA receiver filter bank removed for modifications.Recvr under operational status except for the filter bank.
ansanolla@naic.eduMonday, April 25, 2005 14:50:59ALFA filter bank on system again.Modification that was performed is the relocation of the 11/14 db gain amps for the 1665 mhz side only.
ansanolla@naic.eduTuesday, April 26, 2005 13:06:38There it was a problem with the remote filter bank control and was that channel B of the 1420 Mhz side had the switching trasistor socket shorted out.Also switching time looks to be too short.
ansanolla@naic.eduWednesday, April 27, 2005 14:09:08Cryo lines swapped by JV to the original ones,11(R)& 12(S).Lines in use were 19(R) & 20(S).
ansanolla@naic.eduMonday, May 9, 2005 13:13:37Cryo comp been used in ALFA 2 was swap from ED 009 to ED 011.The cryo comp ED 009 belongs to the SBH receiver.
ansanolla@naic.eduThursday, May 12, 2005 10:13:17On 5/9/2005 cryocomp ED 011 was installed as ALFA cryo comp #2 . It was been using cryo comp ED 009 that belongs to the SBH , it was replaced because the SBH was to going to be used on 5/11/2005.But yesterday the cryo comp ED 011 (ALFA #2) devoleped problems , so the only alternative at the moment was to use the 430 Mhz cyocomp ED 007 while preparing the spare one (not set for the gimble).
ansanolla@naic.eduThursday, May 12, 2005 15:21:05Cryo comp ED 012 installed on ALFA #2 cryo system , press were 80/280 psi.
ansanolla@naic.eduMonday, May 23, 2005 15:36:35SB interlock sw installed in the ALFA recvr.
ansanolla@naic.eduThursday, May 26, 2005 14:59:44ALFA interlock sw was relocalized by JVR. It was not making contac.
cheojrosa@naic.eduWednesday, June 22, 2005 08:57:00on 8 june 05 frig #2 locked up, 13 june turned off both compressors. On 15 june coco replaced the two frigs. we lost 100psi in one day in compressor two. yesterday we discovered the cause ( pure luck). during a power transfer ( from house to commercial power) the relief valve popped on frig number 2. coco thinks its the compressor thats causing it.
W. Genterwgenter@naic.eduTuesday, June 28, 2005 10:03:02Following the crosshead replacement on the 15th of June 2005, the first stage amplifier on beam 1 channel b was seen to draw to much current. Bias cables were switched with the next connector, bias cards were interchanged. This did not repair the problem with this amplifier. Measurements were made with Ganesh by adjusting the Vd and the conclusion is that there is a resistance in parallel with the Vd for the Amplifier inside the Dewar. The most likely cause is icing which resulted from the last cool down. We are looking for time to make warm the Dewar and re-cool following pulling a vacuum of less than 1uc. This will take place as soon as we are able to have the time.
ansanolla@naic.eduTuesday, July 26, 2005 13:05:08Found the 5&6 allen wrenches.Were drop from service platform.`
W. Genterwgenter@naic.eduMonday, August 8, 2005 06:37:05Alfa shut down for warm-up at 0030 this morning.
W. Genterwgenter@naic.eduWednesday, August 10, 2005 09:34:55When ALFA is warm the stage readings are as follows: K3 70K = 262.8K K4 20K = 303.4K 70K = 248.3K
ansanolla@naic.eduThursday, August 11, 2005 14:44:17Yesterday before 430 mhz dual beam run we tried to start the cool down of the receiver who has been pumped since Sunday. But the vacuum at it was higher than 9 un's , so we aborted the cool down ,til Friday.
ansanolla@naic.eduFriday, August 12, 2005 14:40:36ALFA receiver cry system turn on.Its vacuum was 6 un's , So after turning cryo system on vacuum lowered to 3.5 un's. Cryo comp ED015 (ALFA 1) was charged to 280 , it was 20/250 , now is 40/280.
ansanolla@naic.eduFriday, August 26, 2005 13:02:57Due to power failure on 8/25/2005 system warmed up.Cooled.
ansanolla@naic.eduThursday, September 15, 2005 13:11:05The ALFA Dewar been pumped in both ports.Its cryopump drivers were swap to external ones while LM & JR were working with the power cable in order to get rid of the 327 Mhz rfi.After they finish the their installation internal cryopumps driver back.Also the Dewar drive motor shut off till tomorrow.
ansanolla@naic.eduThursday, September 15, 2005 16:04:20Due to wrong power phasing on the Hoffman box that control Dewar rotation/Crosshead driver , the system was warming up.Both cryoheads were running backwards.It happened when LM & JR rewired the power cable into the box in order to shield it to avoid the 327 Mhz rfi.Swap phase at the power cable and walla!!!! , back to normal.
ansanolla@naic.eduMonday, November 7, 2005 14:34:06ALFA Dewar at room temp , been pumped.Start cool down on thursday.
ansanolla@naic.eduWednesday, November 9, 2005 11:21:11The cool down was delay due to contaminated vacuum pumps , the Dewar vacuum was above a 100 un's.3 pumps on it.
ansanolla@naic.eduFriday, November 11, 2005 11:07:21System been cooled.Also Filter Bank back after modifications (new filters on it).
ansanolla@naic.eduTuesday, November 15, 2005 09:23:56WG came to close valves on Dewar after its cool down on Sunday. Also filter bank was checked and is OK ( new 1440 Mhz filters).
ansanolla@naic.eduWednesday, November 30, 2005 13:38:42A 6 db pad was added to the downstairs amps from channel 0 to 5 in both channels A & B. At channel 6 of both channels a 10 db pad was added.All attenuators will be removed on Friday.Any question look for Bill Genter.
ansanolla@naic.eduFriday, December 2, 2005 09:52:06The pads at the downstairs ALFA amps were removed , were 6 db from channel 0 to 5 A & B. Channel 6 A & B were 10 db pad.
ansanolla@naic.eduFriday, January 13, 2006 15:58:45On 12/15/2005 was installed the ALFA IF attenuator sw unit.
ansanolla@naic.eduTuesday, January 24, 2006 15:41:45ALFA dewar warmed up.Amp #6 of channel A failed.Been pumped.
ansanolla@naic.eduWednesday, January 25, 2006 10:49:26ALFA Dewar vacuum is 3.5 un's.Amp #6 on chann A still bad after warm up.
ansanolla@naic.eduThursday, January 26, 2006 11:50:39ALFA amp #6 of channel A problem comes inside the Dewar.Cool down started at 11:00 A.M.
ansanolla@naic.eduFriday, January 27, 2006 15:12:31System back to service.Amp # 6 on channel B alive.
ansanolla@naic.eduMonday, February 13, 2006 13:03:23System down for the replacement of crosshead ED 021.Turn off on 2/10/2006.Cryohead replaced today.
ansanolla@naic.eduWednesday, February 15, 2006 15:59:43ALFA back into action.Vacuum at the pump was .8 un's and at the Dewar .6 un's.Both valves closed and called to CR to tell them that ALFA recvr was ready.
ansanolla@naic.eduFriday, February 17, 2006 15:27:08ALFA cover removed.
ansanolla@naic.eduThursday, March 9, 2006 13:08:08JV modified the Alfa shutter , so that a single person can handle it.He redesing the holding braket so that it can hold into place alone.
ansanolla@naic.eduTuesday, May 9, 2006 14:40:42ALFA cryo comp ED 015 was replaced by ED 012.ED 015 looks to have a bad flapper due to high preassure in the low side (150 PSI), but was not the compressor it was the cryo head , that was replaced.
ansanolla@naic.eduFriday, June 2, 2006 11:33:03ALFA receiver been used by the SERENDIP people.They will be using the 14 channels of it.That means 0/6 A and 0/6 B at the same time.
ansanolla@naic.eduMonday, June 12, 2006 12:48:00ALFA cryo comp 2 (ED012) has a faulty flapper sound , notified.
ansanolla@naic.eduMonday, June 12, 2006 12:52:33ALFA faulty cryo pump # is 1 , not 2 as I wrote.
ansanolla@naic.eduTuesday, August 15, 2006 13:52:50ALFA 2 cryo system was replaced by the SBH system due to a leaky line in the receiver cable wrap at SB xmitter room.So SBH system is hot.
ansanolla@naic.eduThursday, August 17, 2006 09:06:38System will be running with the SBH system until new cryo hoses arrive.So SBH will be down for a while.
wgenterwgenter@naic.eduMonday, August 21, 2006 17:12:56This afternoon ALFA beam 1 CH B indicated a high current and 0 VD. First we isolated the problem as external to the DEWAR. Next we changed the bias card (F98). Still we had a problem. Next we changed the MUX card, now the amps are working well in beam 1 CH B.
ansanolla@naic.eduFriday, September 1, 2006 11:18:19The ALFA 2 cryo comp was hooked to cryocomp ED 009(is SBH cryo comp)last week due to a leaky hose in the cable wrap. Hose was replaced yesterday by CR & JVR.Today CR put both systems back to normal, SBH recvr with ED 009 and ALFA 2 with ED 014.SBH being cooled.
ansanolla@naic.eduTuesday, September 5, 2006 14:31:24ALFA 2 still leaking , supply was 210 psi this morning.After looking for the leak CR charged to 75/290.Didn't localized the leak , will work on it tomorrow.
ansanolla@naic.eduWednesday, September 6, 2006 16:21:22System cryo is not leaking.
ans anolla@naic.eduThursday, September 7, 2006 11:31:27Channel 5B was checked for unstability , can't duplicate. It was reported on Sunday.
ansanolla@naic.eduTuesday, September 26, 2006 10:51:54Cryo system still leaking.Found it at 220 psi.
ansanolla@naic.eduThursday, September 28, 2006 13:17:58There are problems with the cryogenic system 2. There is a He leak of around 10 psi daily .Vives & Carlitos are working with it , they are trying to pint point it.So yesterday they replace return hose # 11 with hose # 20 (spare) , but today they found that there is no leak with hose # 11.So they removed today line # 12 (supply) an replaced it with # 20 , in order to isolate the problem. During these last hose swap they found that cryo pump on these channel quit working properly.
ansanolla@naic.eduMonday, October 2, 2006 12:00:37ALFA down for crosshead replacement on ALFA #2.
ansanolla@naic.eduWednesday, October 4, 2006 13:14:15ALFA being cooled.
ansanolla@naic.eduThursday, October 5, 2006 13:00:12Yesterday when starting the cool down it was found that the replacement cryo head was leaking.So after pinpointing the problem that was that the "O" ring between the motor and the Scotch yoke assembly was missing , Ha! Ha! Ha!_____Ha!.Somebody?????? forgot that it needs that "O" in order to avoid a He leak.
ansanolla@naic.eduFriday, November 17, 2006 11:20:42ALFA being warmed.
ansanolla@naic.eduTuesday, November 21, 2006 14:40:59Vac valves were closed.Big pump was 1 un , small 2 un's.Pumps removed from dewar.
ansanolla@naic.eduWednesday, November 22, 2006 16:09:13ALFA back to service.
ansanolla@naic.eduMonday, April 16, 2007 10:33:38ALFA turned off on 4/11/2007.Being disasamble to be taken down on 4/17/2007.
ansanolla@naic.eduWednesday, April 25, 2007 13:10:42System cryo comp taken down for service.
W Genterwgenter@naic.eduMonday, June 11, 2007 09:51:55Mux Relay Card F36. All relays were replaced by J Rios.
CRVThursday, January 3, 2008 09:45:08Alfa back on the Dome on Dic.10th,2007. Vacuum pumps were connected and pumping.
CRVThursday, January 3, 2008 09:50:44Alfa cooldown Startedon Dic.18th,2007 with two external Cross head driver because the control box isn't ready.
CRVThursday, January 3, 2008 09:51:50Alfa control box lifted to Dome on Dic 19th,2007
CRVThursday, January 3, 2008 10:26:16Alfa cooling stopped on Dic.21th,2007
CRVThursday, January 3, 2008 10:27:15Alfa cooling restarted on Dic.24th,2007 at 14:20
CRVThursday, January 3, 2008 10:28:31Alfa control box connected By Jorge an Willy Iguina on Dic.27th,2007.
CRVThursday, January 3, 2008 10:30:16ALfa was reconnected to the crosshead driver on the control box on Dic. 27th,2007
crvThursday, January 3, 2008 10:36:05Alfa filter bank problemwith chanel 4B. We used relay S/N 2665 today to replace the bad one with S/N 1721 in channel 4B of the ALFA filter bank 100 MHz path.
Dana Whitlowdwhitlow@naic.eduMonday, February 18, 2008 10:02:47Added relay coil "catch diodes" to driver chassis for ALFA 100-MHz filter bank selection. Without the diodes, the peak voltages on the driver hexfets was significantly in excess of device ratings when the driven coil was turned off.
crvThursday, April 17, 2008 16:15:49Bias card F98 on Chanel 5 was replace due to problems on chanel 5a amp.
crvMonday, May 5, 2008 10:37:27ALFA was warmed during the and will be cooled during the 430MHz transmitter run.
crvWednesday, May 7, 2008 09:19:52Tuesday May 6th the dewar vacuum valves were opened and vacuum started at 6:40 am.
crvWednesday, May 7, 2008 09:23:15Today alfa started cooldown at aprox. 6:45am. Dewar vacuum was at 5 microns. The small Vacuum pump valve was closed and the pump left on .
crvFriday, May 9, 2008 09:44:31Alfa vacuum valve close and pump removed. Alfa rotation tested and working. Alfa amp. cxhecked and working no fault indications. Alfa's compressor charged normal after a cooldown.
Dana Whitlowdwhitlow@naic.eduTuesday, September 23, 2008 17:21:15ALFA MODS 19 Sept 2008 On all 14 channels, we performed the following mods to the ALFA bias cards (changes to stage 3 only): a) Removed the Vdg clamp zeners (6.2V) on stage 3 only. D15 (pol A) and D39 (pol B). b) Modified the stage-3 Vd reference to extend the range to approx 6Vdc. This involved removing the jumpers JP9 & JP10 (pol A) and JP21 & JP22 (pol B), then adding 12k resistors from +15V to the tops of VR5 (pol A) and VR11 (pol B). c) Adding two 2.7k pulldowns to the +5V reference supply line on each card, to swallow the current supplied through the added 12k resistors in (b). This was done to ensure that the load current on the +5V reference supplies exceeded the minimum required load for the voltage regulator IC's being used. d) Adjusted all stage-3 biases to 5.9V and 17 mA. Note that these changes pose the risk of slightly degraded drain bias voltage stabilities on stage 3, depending on much noise and/or drift occurs on the +15V supply. These were performed in the expectation of reducing the severity of the lengthy radar overload recovery problems that have been seen.
Stephen Jensensjensen@naic.eduMonday, August 3, 2009 16:52:40Power failure last night. Only one crosshead is working on ALFA. The other crosshead makes a loud sound & the piston may not be working. The amplifiers are cooled down to 30 Kelvin.
crvThursday, August 6, 2009 15:16:30reciever started vacuumed today 8/6/2009 after sunday night power failure and wednesday power failure .
crvMonday, August 10, 2009 10:09:29system vac. was restarted on friday august 7th 2009 because the vacuum pumps where full of water extracted from the dewar. Vac. pump where clenead and repaired and vac. started at 4:00pm
crvMonday, August 10, 2009 10:12:16Dewar vac. was 3 microns cooling started today at 8:30 AM 20K2 temp sensor bad due to the water on the dewar. Temp. board tested and is good.
Dana Whitlowdwhitlow@naic.eduFriday, November 13, 2009 15:15:34September 13, 2009: ALFA taken out of service in preparation for bringing down for maintenance.
Dana Whitlowdwhitlow@naic.eduFriday, November 13, 2009 15:17:30October 30, 2009: ALFA raised back to platform and bolted in place.
Dana Whitlowdwhitlow@naic.eduFriday, November 13, 2009 15:19:45November 12, 2009: First light (so to speak) for LNA LEDs. Bias conditions not yet optimized for best behaviour.
Dana Whitlowdwhitlow@naic.eduFriday, November 13, 2009 15:21:05November 13, 2009: First pass at optimizing LNA bias conditions for use with LNA LEDs turned ON. Now we wait...
crvTuesday, December 22, 2009 14:59:24ALFA cryo system was turned off on Tuesday December 1st 2009.
crvTuesday, December 22, 2009 15:07:06On Friday December the 4th alfa compressors ed010 and ed016 where replaced for two new scroll compressor(C1, C2). The lines were purged and connected to a test dewar. No signs of contamination.
crvTuesday, December 22, 2009 15:10:51 Monday December 7th 2009 both cryo pump where replaced for two new ones. The new cryo pump are ed040 , ed041. cooldown started that same day.
crvMonday, January 4, 2010 16:51:26ALFA'S C2 compressor has now three cryo pumps (alfa 2, 327MHz , XB).XB ans 327MHz systems were removed from Comp. ED015. These recievers were connected to C2 to use the compressor at maximun capacity and comp.ED15 for SBH.
crvWednesday, January 13, 2010 10:12:25The system is warming up due to a leak on C1 compressor lines. leak will be totally repaired on 1/13/10
crvcrios@naic.eduTuesday, January 19, 2010 13:43:25The cryo system had a damaged o-ring on the connection of the rotary floor. O-ring was replaced on 1-14-2010
criosMonday, February 22, 2010 14:43:30Alfa compressor C1 has the 327 MHz and XB connected and compressor C2 has SBh.
criosFriday, April 30, 2010 17:28:42On April 23rd comp. C2 failed. Alfa's cryo pump 2 was changed to Comp. C1.
crvFriday, October 1, 2010 13:17:59alfa wasturned off warming and doing vacuum.
crvcrios@naic.eduTuesday, October 5, 2010 15:50:42 Alfa started cooling down today at 6:30am
crvThursday, October 7, 2010 12:00:40Amp 2B and 6B had bias settings problem. 2B was adjusted and works and 6B needs more testing.
crvThursday, October 7, 2010 12:55:33Amp. 6b was checked and the open is inside the dewar. The system will be warmed to see if we can eliminate the open.
crvMonday, October 11, 2010 15:41:55system started cool down on friday October 8th 2010.
CRVMonday, October 11, 2010 15:44:08 Ganesh cane to check Beam 5 on sturday October 9th 2010 because one amp. didn't have any Vd and Id. He replaced the bias and mux card.
CRVMonday, October 11, 2010 15:45:28System wasn't totally cooled we re-charge the compressor to 280 PSI. after 30 min temps. were normal.
crvMonday, October 11, 2010 16:10:57Vd voltages on both channels of beam 5 were adjusted from 5.44v to 5.92v
crvFriday, October 15, 2010 11:45:24The bias board of beam 5 was change for the original board. Bias settings where set to normal values.
crvcrv@naic.eduMonday, October 18, 2010 16:55:37The system had a helium leak and it was traced to a hose that connect the floor lines to the Alfa snake lines. The hose was replaced and system is still cold.
Dana Whitlowdwhitlow@naic.eduThursday, December 9, 2010 14:40:59In response to complaints of unstable gain and bias in ALFA beam 5 LNAs, brought down bias card and tested it in lab. It appeared to be very stable, so reinstalled it in ALFA and looked for bias oscillations. Found same at ~2.5 MHz, was able to quench the oscillations bt increasing bias current in CHB to approz 19 mA. We'll see how long this lasts.
crvcrios@naic.eduMonday, January 31, 2011 14:46:53The system was charged with helium the compressor was al 200 psi leak is been checked.
crvMonday, February 7, 2011 14:03:37The system had a mayor leak and the helium lines 11 and 12 were swaped with the spare lines 13 and 14.
crvMonday, February 14, 2011 14:18:25The system still has a small leak. The two T connection were replaced for two new connectors. system was recharged.
Dana Whitlowdwhitlow@naic.eduWednesday, February 16, 2011 13:50:49Open-loop bias card which has been running in beam 2 (for the last year or so) was moved to beam 1 today during diagnostic tests of problem with a small (0.5% Tsys) pulse in receiver output power seen ~every 5 sec. Open-loop card was left in beam 1 slot following completion of tests. Problem was only just noticed, but review of past records suggests that it has been happening for at least several weeks if not months. Problem was found to be associated with scanning by the bias sampling system for remote monitoring and is much more severe with the open-loop bias configuration. Full solution is TBD, but a likely initial response will be re-configuring the bias card for closed-loop operation, since the open-loop configuration never did really solve anything.
crioscrios@naic.eduWednesday, March 16, 2011 14:49:34The dewar cryo pump was turned off on monday 3/14/2011 for schedule maintnance. Tuesday 3/15/2011 The cryo pump were replaced and the dewar vac. was on 30 microns. Today the vac. was on 3.5 microns and the cryo pump were started. System is cooling.
Dana Whitlowdwhitlow@naic.eduMonday, May 2, 2011 17:33:552 May 2011: Made adjustments to ALFA for the usual reason of resolving gain instabilities. It was found that beam 5 was oscillating strongly at just a few MHz, but both 5A & 5B had to be on simultaneously for this to occur. Tried various things and found that further reducing Vd on stage 3B was the trick. However, could not kill oscillation entirely, just reduce it by a fair amount over a fairly narrow range around 5.32V. This worked best for drain current around 17 mA. At this "minimum" condition the oscillation level was very erratic, varying randomly over a range of at least 10 dB. Then discovered that turning beams 4 back on nicely killed the oscillation altogether, and it stayed dead even with all the non-6 beams turned on simultaneously. Next we took a look at gain stability on selected beams, using the VNA function of the R&S ZVL instrument and injecting the signal into the port on the CAL box. Overall S21 of this path (from injection port to dewar RF output) was approx -40 dB. With the VNA IF BW set to 100 Hz and injected input levels of roughly -10 to 0 dBm, the VNA noise fluctuation level was roughly 0.01 dB and readings at 0.1 dB/div were entirely practical. All the ALFA amps tested this way (beams 5A, 5B, 4A, 4B, and 3A) showed similar gain variations of up to ~0.1 dB over a time span of a few minutes. To be sure that this was not due to the VNA, a check was also performed with the VNA ports connected straight through, and no such variations were seen even when the VNA power was set so low that the normal VNA noise level fluctuations rose to higher levels than were seen during the tests on the actual LNAs. Following these measurements we rechecked for oscillations and found none.
CRVcrios@naic.eduFriday, May 6, 2011 15:29:52On the RF to fiber converter box of ALFA we swaped the board of beam 6 with the board of beam 1. We also reinserted each of the others beam board.
Denis Urbaindurbain@naic.eduMonday, May 9, 2011 11:28:43problem with Tsys on ALFA over the weekend was due to power strip that was left on the OFF position. Problem resolved by turning the power strip back on
philphil@naic.eduMonday, May 23, 2011 11:28:43ganesh readjusted bias voltages for beam 1 back to 16feb11 values. The 16feb11 values gave the best stability when the leds were used with the amplifiers
Dana Whitlowdwhitlow@naic.eduFriday, June 24, 2011 15:48:04On Thursday June 23, 2011, the bias card for beam 2 was replaced in an effort to reduce instability noted on 2A. The problem immediately addressed was that 2A showed about 2x to 3x the output noise variation as did 2B on a time scale of seconds. The bias card replacement seemed to help. However, the high noise temp on 2A was not improved, and the diurnal variations in stage 1 bias current has not yet been checked to see what the impact of the bias card change was.
Dana Whitlowdwhitlow@naic.eduSunday, July 17, 2011 16:18:09Repaired the F98 bias card removed from service in June. The fault lies with the Vd-Vg clamp zener having gone bad, acting like its clamp voltage has gone way down. Not having a spare on hand, I simply moved a jumper from JP4 to JP3, tantamount to substituting a 5.6V zener for the more proper 4.7 V zener.
Dana Whitlowdwhitlow@naic.eduSunday, July 17, 2011 16:21:02Today Ganesh & I replaced the MUX board from Beam 4, to correct a problem wherein beam 4B could no longer be turned on. While working with this we discovered the surprising and yet-unexplained fact that the beam could be turned on in local control mode, but not remotely.
Dana W. & Denis Urbaindwhitlow@naic.eduFriday, December 2, 2011 17:37:39Today (02Dec2011) we replaced all the home-kluged bias filter cables with the nice shiny commercially-built filters recently acquired. No oscillations were observed, and all looks fine- EXCEPT that the LEDs on beam 2A quit working. This problem has happened before, according to Denis, and it fixed itself, so we did not spend a lot of time troubleshooting this LED failure as we had much other work to accomplish in the dome.
crioscrios@naic.eduThursday, February 2, 2012 10:04:31A He leak on the system was detected on Jan 24th 2012. The helium line 11 was replaced with alfa Spare J. The leek on line 11 will need welding on a copper connection.
Dana Whitlowdwhitlow@naic.eduFriday, May 4, 2012 14:08:11BRIEF REPORT ON AO RADAR BLANKER PROBLEM SEEN 03/04MAY2012 LBW observer complained that the radar blanker was not doing its job, and it seemed not to be locking. First obvious symptoms noted: the unlock light was flashing at a regular rate (~2 Hz?), the VCO tuning voltage was indicating at or near the high end of its range, and the "acquired signal level" was reading near zero. A quick check with a 'scope revealed that the IF signal looked very healthy and that the 5-wide radar pulse sequence looked normal, but the blanking pulse train was a bit late and was sort of oscillating over time positions ranging from a tiny bit late to several hundred microseconds late. The oscillation period was somewhat variable, but was typically a modest fraction of one second. I captured ~10 sec of data from the "detected output" port of the radar blanker with a DPO4104 with 10,000,000 record length and did a Fourier analysis, and found a sequence repetition rate of 355.72 Hz. But then, unbeknownst to me, Phil called the radar site to inquire, and checking revealed a COHO lock loss at the radar. They relocked it, and our radar blanker began working properly again. I redid the 'scope measurement and now found a sequence repetition rate of 354.53 Hz, which is about 0.36% lower. So next time this happens, check the SRR and if it's not well within 0.3% of 354.528Hz (probably should be well within 0.1%), blame the radar first. Dana Whitlow 04 May 2012
Denis Urbaindurbain@naic.eduThursday, November 15, 2012 13:11:35The cooling circuit of the receiver was losing Helium. The leak was small initially but became a problem when we had to recharge the system everyday. The loss in Helium pressure reduces the cooling capacity of the system and was noticeable on the receiver tempertures. The leak was found in the cable wrap and the line that was leaking was replaced.
Dana Whitlowdwhitlow@naic.eduSunday, January 13, 2013 08:41:32Yesterday, Sat 12Jan 2013, ALFA cooling was shut down and vacuum pumping was initiated in preparation for cryogenics maintenance which will continue on the coming Tuesday morning. Also the ALFA cover was left attached, the amplifiers were turned OFF, and ALFA rotation was disabled (due to the vacuum pump's being connected).
Denis Urbaindurbain@naic.eduWednesday, January 23, 2013 16:41:56 The ALFA receiver had been running for more than 6 months, we decided to preventively warm-up the receiver to clean the cold heads and restore a good vacuum during a long dual beam run. After the warm-up, some of the LED on the amplifier stopped working and 5 of the temperatures sensors were not giving readings. Some water was found on the cold head displacers proving the need for service. The vacuum pumps connected to the Dewar had to be cleaned because they were contaminated with water. This shows that the Dewar had accumulated a lot of water over the several month of service and it was time to restore the vacuum. Unfortunately the thermal cycle had a negative impact on the LED of the Beam 3B and 5B. The problem with the temperature sensors is worrisome because 5 out 8 stopped working; it seems that they are shorted. Is the problem coming from the sensor or the cable, we don't know.
Dana Whitlowdwhitlow@naic.eduFriday, September 27, 2013 14:44:52Last night, Thurs 26 Sept 2013, the RF switches for 5B were replaced in the ALFA filter chassis. The previous user had reported that 5B was essentially dead in the unfiltered position, but apparently OK in the filtered position. Thanks to Denis, Guarionex, and Felix for their hard work in the dome under intense pressure.
Dana Whitlowdwhitlow@naic.eduFriday, November 1, 2013 14:43:28This morning we looked for the cause of the reported severe instabilities in ALFA beam 2A. First we looked by monitoring the level of a CW signal injected from the primary reflector surface with a small helix antenna and sig gen, in the hopes of being able to detect minute changes in level on a very short time frame. We examined the signal at the output of the ALFA filterbank chassis, before 'Eddie's box', and we stimulated ALFA by wiggling cables and banging on the dewar with the usual 2x4. No effects were seen other than slow and pretty smooth variations in signal strength attributed to path variations. In order to confirm this last conjecture, we tuned away from the CW signal and looked at system noise at the same point, in a 3MHz bandwidth. Here the useful measurement sensitivity was reduced, but we could easily have distinguished misbehavior at the 0.1dB (2.3%) level. With this method we saw no overall instability (other than two faint sky sources that drifted by), and no response to cable wiggling and pounding on the dewar. Here we also tried wiggling the trimpots setting the LNA bias values on beam 2A. There was no visible response to Vd, but Id changes were visible as small but smooth variations in system noise level as the 10-turn trimpots were rocked back and forth over about a 1/2 turn p-p range. There was no evidence of erratic pot slider contact behavior. In a subsequent conversation with Phil I learned that he had already searched for correlations between bias value changes and the reported instances of beam 2A instabilties, and found none. So the problem reamains unsolved at this point.
Dana Whitlowdwhitlow@naic.eduWednesday, November 13, 2013 14:58:52Yesterday we continued the search for the cause of the severe instabilities repeatedly reported on beam 2A. This time the subject of inquiry was part of the signal path beyond the dewar and filter box. We passed the downstairs 260MHz IF for both pols of beam 2 through 40MHz-wide BPF's, distribution amplifiers, into squarelaw detectors. Then we beat on the downconverter (aka "Eddie's box") using the UPS-traceable whacking tool (an 8-inch length of 2x4), we wiggled and tugged on the fibers where they exited the rear of Eddie's box, etc. In no case was there any clear correlation between any of these acts and misbehaviors of 2A. However, during the fiber tugging exercise there were some mild disturbances to both pols, but the effect was not repeatable. In general, I could comment that 2A seemed slightly more unstable than 2B. The results of this were really not conclusive, and I plan a near-future test in which we remove the front panel of Eddie's box and "stimulate" individual fibers right at their individual connectors.
Dana Whitlowdwhitlow@naic.eduMonday, November 18, 2013 16:17:56This morning we pursued the beam 2A instability problem further, beginning by watching 2A with a squarelaw detector while Felix et al wiggled and tugged on fibers where they connected inside Eddie's box. There was no obvious correlated response, but we did see (for the first time) numerous seemingly random occurrences whose nature correspond closely with the description of the problems repeatedly reported by R. Minchin. These were generally step changes (or perhaps better described roughly as rectangular pulses) of 5-10% of Tsys in magnitude, with durations ranging from a bit less than a second to minutes. Sort of like extremely slow "popcorn noise". We then began looking into possible causes in the WAPP room, concentrating on the fibers coming in from upstairs, the fiber receivers, and the switchable attenuators in the coax environment, but were unable to elicit any effects resembling the "prime symptoms". We did discover, however, that certain bending/twisting motions on the fibers could induce visible transmission variations, especially if done in the vicinity of the sharper bends where the fiber bundle is hanging from something. But I believe this to be a red herring. As a trial, we interchanged fibers 5 & 6 at both ends (so the pols were not crossed), and watched. The problem clearly stayed with Pol A, so we can probably forget about the fibers themselves. Knowing that today we were actually seeing the prime problem, we went upstairs after lunch and repeated the RF tests looking with a SA, both at the dewar output and at the filter chassis output, and did not see the problem at all. But we clearly were seeing it immediately before leaving for the platform. NEXT IDEA: use CW signal injected at various points in the system to see if we can see and localize the gain jumps. We should probably begin downstairs for convenience, initially inserting the signal just downstream from the fiber receivers in the WAPP room, and watching through the usual squarelaw detectors.
Dana Whitlowdwhitlow@naic.eduThursday, November 21, 2013 16:15:36Today we conducted further investigations into the beam 2A severe instability. Yesterday, Felix had replaced the beam 2 card in Eddie's box and also undid the test fiber swap at both ends. At no point in this morning's tests did we see any evidence of the big problem. We also temporarily went back to the old card, still no evidence of the problem. However, with the old card in place, wiggling the card in its socket did cause substantial changes in level for both pols, with pol B being the worst. With the new card put back in, the wiggling effect was much smaller, so we finished the day with the new card left in operation.
Dana Whitlowdwhitlow@naic.eduThursday, November 21, 2013 16:19:59Yesterday, Felix replaced the beam 2 card in Eddie's box with a spare, and verified that Tsys was OK. This morning we looked into the effects of the new card. There was no evidence of the instability problem. We then briefly put back in the old card, still no problem. Then we wiggled the card in the socket, yielding problem-like variations in both pols, but somewhat less severe than the real problem. With the new card put back in, the wiggling sensitivity was greatly reduced, so we finished the day by leaving the new card in operation.
CAL CTRL PROBLEM on remotedwhitlow@naic.eduWednesday, February 26, 2014 10:07:03Experienced a problem wherein the downstairs-originated TTL CAL ON/OFF signal would not control the CAL state when the upstairs control unit was set to 'remote', which is its normal operational setting. Also, the CAL status indicator near the bottom right of the ALFA control window said "ILLEGAL". We finally discovered that changing the CAL LEVEL to its other state, then back to the desired state, fixes the problem. We do not yet understand what was going on- that will presumably be the subject of a future entry on this topic.
ALFA Cold-head servicing and Helium leaksfelix.fernandez@naic.eduTuesday, March 18, 2014 07:52:00ALFA's cold-head #2 had been having slight "knocking" issues since FEB-15-2014. Here's the repair progress and current status. - MAR-10-2014: ALFAs's cold-heads had reached the end of their operational life. Thus, ALFA started the warm-up process on the same day. - MAR-12-2014: Removed both of ALFA's cold-heads and started the vacuum process using two pumps. - MAR-13-2014:Prepared two cold-heads to replace ALFA's current ones. - MAR-14-2014:Replaced both ALFA's cold-heads and continue vacuum process. - MAR-15-2014: Excessive bias currents on beams 3A and 5A were reported. After debugging, it that the problem lays inside ALFA. On the same day the compressor was found to almost out of Helium. Therefore, the compressor was recharged. - MAR-16-2014: The compressor had leaked almost all the Helium that was inserted the day before. It was concluded that there is a significant Helium leak on the system and that the system has bee compromised by external (air or humidity) contamination. - MAR-17-2014: After checking all the lines the only visible leaks was a fairly small one on the cold-head's (#2) relif valve. Further inspection showed that there were two leak inside the compressor. One leaks was detected on a stainless steel line and the second one on the compressor's absorber. Both the stainless steel line and absorber were replaced.
Starting cool-down process felix.fernandez@naic.eduTuesday, March 18, 2014 15:00:59The compressor, helium lines, and cold-head were purged and pressurized. The cool-down process started at 14:30 with the compressor supply pressure reading 310 psi, and a vacuum reading of 6.5e-3 Torr on ALFA's vacuum monitor.
Cooldown finishedcrios@naic.eduThursday, March 20, 2014 15:22:47Alfa finished the cooldown and the vacuum pumps were removed. The receiver is ready to be used.
ALFA down for maintanacefelix.fernandez@naic.eduTuesday, April 8, 2014 09:38:26ALFA was brought down for maintenance on MAR-25-2014. The main reason for this decision was the fact that beams 2A, 3A, and 5A were unusable after the last thermal cycle. Beam 2A was unresponsive to cal signals or "birdies". Beams 3A and 5A both had bias problems in their second stage as VG was unresponsive to the feedback signal. Furthermore, all temperature sensor were offline. The expected return date for ALFA is MAY-7-2014.
Dana Whitlowdwhitlow@naic.eduTuesday, May 27, 2014 10:51:49ALFA has been re-installed and calibrated. All beams and pols are working and pointing is OK, per tests run by Phil Perillat.
Compressor's oil leakfelix.fernandez@naic.eduThursday, July 10, 2014 15:38:12An oil leak has been found in ALFA's compressor. The leak's rate is rather small and no He pressure has been reported. The oil leak seems to be coming from the cool-oil's hose between the heat exchanger and the scroll compressor. We will repair the compressor next week after this weekend's observations.
Beam 6B - Stage 1felix.fernandez@naic.eduWednesday, October 8, 2014 15:09:58Tsys values of Beam 6B were suspicious. After debugging we concluded that Beam 6B's amplifier's first stage is non operational. We checked the bias cable, bias card, analog MUX, analog relay and all components outside the Dewar. The only possible explanation is that there is either a faulty connection inside the Dewar or that the Gate of Beam 6B - Stage 1 has been damaged. We will keep Beam 6B OFF until ALFA is brought down for repairs.
Single Cold-head Modefelix.fernandez@naic.eduMonday, October 20, 2014 11:44:44ALFA's Cold-head #2 had a mechanical problem on Oct-17-2014 morning. The cold-head stopped working (the motor seized) during that time frame. ALFA has been working with only one cold-head since then. We must warm-up ALFA completely so that we may service the cold-head.
ALFA thermal cycle - STARTfelix.fernandez@naic.eduTuesday, October 28, 2014 15:44:02A thermal cycle was started on OCT-27 to service ALFA's cold-head #2. After removing both cold-heads there was humidity found on both of them. Cold-head #2 had a broken rear motor bearing. This bearing was replaced and the displacers were baked. Cool-down should start later this week.
Cold-hear ratchetingfelix.fernandez@naic.eduWednesday, December 24, 2014 10:38:52On the week of DEC-8-2014 cold-head #2 was ratcheting and a thermal cycle was started to enable cold-head servicing. During a week long T1193 we found a minor mishap in the vacuum system. Namely, one of the braces came off and vacuum was lost. Hence, we restarted vacuum on DEC-18-2014 and cool-down started on DEC-22-2014. Finally, ALFA is back online today DEC-24-2014. Merry Christmas!
ALFA's trip to the cryo-labfelix.fernandez@naic.eduFriday, May 8, 2015 09:41:14On MAR-12-2015 ALFA was brought down for repairs. The main issues were: B5B and B6B were not biased properly although the bias cards seemed to be functional. After a detailed debugging both problems were a combination of faulty analog MUX cards and fault bias cables. During mainteance we completed the following tasks: (1) Fixed a faulty F36 analog MUX card by replacing a bad NPN Darlington (2) The spare F36 had a similar issue but this time it was a faulty analog MUX that was replaced (3) Fixed two spare F98 bias cards by replacing all 4.7V Zener diodes. (4) Fixed a faulty bias cables at the Dewar interface by re-soldering the wires and replacing all connectors with 90deg. bent connector, this to reduce cable strain. (5) 70K1 had a lose solder joint inside the dewar that was repaired. After all these repairs ALFA was brought up on APR-28 NOTE: ALFA's remote vacuum sensing is operational but for some reason the CIMA interface has the wrong scale.
Cryo pump replacementcrios@naic.eduThursday, July 9, 2015 12:52:23Both of ALFA's cryo pumps have been replaced and the system is cooling.
Beam 0B erratic behaviorfelix.fernandez@naic.eduFriday, August 14, 2015 09:51:38R. Minchin reported that beam 0B has some erratic behavior. The system gain drops suddenly and then goes back to its normal value. This behavior was seen during A2048 observation on AUG-02-2015 between 02:00-03:40. Today C. Rios found our that the SMA cable on beam 0B going from the filter bank into the fiber Tx was extremely loose. He re-tighten it in hopes of solving this issue.
ALFA thermal cyclefeix.fernandez@naic.eduWednesday, November 18, 2015 15:06:18ALFA went through a planned thermal cycle last week during HF's first campaing. This was done as a possible fix to bring back to "life" B3B-Stage#1. After the thermal cycle ALFA is back to 100% operation.
Several Issuesfelix.fernandez@naic.eduWednesday, June 14, 2017 11:15:39Beam 5A: currently offline. When the LNA is turned ON the drain current on stage #1 goes to 17mA, Vd = 3V and Vg = -0.618 After debugging we know that Vg is not connected inside the Dewar. Beam 0A: Some erratic behavior has been observed (starting 17-06-10) where Id of stage #1 would ramp to 31mA and Vd would go to 0V. After Turning the LNA OFF for some 24hrs once the LNA was restarted the problem disappeared on its own. 60HZ beat on B5 and B6 both pols: There was a bad Acopian power supply at ALFA's mixer chassis. The power supply was replaced and the problem went away. 120Hz beat on all beams when the narrow band filter was selected: There was another bad Acopian power supply inside the chassis on top of the ALFA's mixer chassis. This power supply provided the 5V rail for the amplifiers placed after the narrow band filter.
Thermal Cyclefelix.fernandez@naic.eduSaturday, September 16, 2017 17:25:08After some issues with beam 0A stage 1 we found out that ALFA's vacuum had increased to 3.94 mTorr. This is typically a signature of frost being developed inside the receiver. A thermal cycle was started on SEP-13-2017 during the DB+ procedure. After reaching an internal temperature of 290K on SEP-15-2017 the vacuum process was started. Once the vacuum reached 8.90 mTorr on SEP-16-2017 the cool down process started. We expect ALFA to reach the nominal temperatures before the ongoing DB observation ends.
Fiber link swap from 5A to 4Afelix.fernandez@naic.eduFriday, March 23, 2018 10:00:04Since we know B5A's LNA is dead inside the dewar we decided to swap B5A's fiber link (downstairs, inside the computer room, next to the WAPS) with B4A. We did this since we know that B4A has been really low in power due to a faulty fiber receiver. For now we will use B5A's fiber receiver to bring down B4A. This is TEMPORARY change. Once ALFA is brought down for repairs we will put all connections back to their original state.
Cold-hed mechanical failurefelix.fernandez@naic.eduMonday, August 6, 2018 13:17:42Cold-head #2 had a mechanical failure and it is now locked up. It will be replaced once a thermal cycle can be scheduled.
Sudden increase in temperaturefelix.fernandez@naic.eduWednesday, May 8, 2019 11:37:13Yesterday 2019-05-07 sometime around 2PM ALFA temperatures started to increase. Once we reached the Dome all cryogenic systems were operational but ALFA's vacuum had increased to 16 Torr. We had to thermal cycle ALFA. System evacuation started today and cool-down should start sometime tomorrow afternoon.
ALFA rotation problemsfelix.fernandez@naic.eduMonday, July 8, 2019 10:21:14On 2019/07/07 J. Hessels reported a problem where ALFA culd not rotate to -91.25deg. It turns out that an electronics chassis from ALFA's top-hat slid off and ended blocking ALFA's rotation past -50deg. Luckily, there were no damages and the chassis was locked into its proper position. ALFA is back online with full rotation as before.
ALFA Serial COM resetfelix.fernandez@naic.eduThursday, August 15, 2019 08:50:39For the past few weeks there had been an issue with ALFA's bias monitoring. Namely, reported bias current would go negative, the bias would not report 0 when turned OFF (we they were actually OFF since Tsys confirmed the OFF state) and there was no control on the vacuum gauge. On 201-08-15 we did a reset (turn OFF/ON) the serial COM on ALFA. This was done in an effort to fix all of the above. The reset seemed to work. We suspect that the battery backup might not be at fault and power glitches might be the root cause.
Sudden Increase in Tempfelix.fernandez@naic.eduTuesday, October 15, 2019 07:58:25On 2019-10-15's night ALFA cryogenic temp started to increase. C. Rios went to the Dome and found that the vacuum level had increased to 7 Torr. A thermal cycle was started then.

Back to Top